In a triangle, one angle measures 70 degrees less than the second angle. The third angle measures 80 more than the first. What are the angles?
Answer by mananth(16946)

In a triangle, one angle measures 70 degrees less than the second angle. The third angle measures 80 more than the first. What are the angles?
...
let second angle be x
first angle = x-70
third angle = x-70+80=x+10
...
add up
x+x-70+x+10=180
3x-60=180
3x=180+60
3x=240
x= 80
..
80, 10, 90
...
m.ananth@hotmail.ca